A 500kg car takes a round turn of radius 50m with a Velocity of 36km/hrs . The centripetal  force is?



1
Expert's answer
2022-06-13T08:23:22-0400

Given:

"v=36\:\rm km/h=10\: m/s"

"m=500\:\rm kg"

"R=50\:\rm m"



"F_c=\frac{mv^2}{R}=\frac{500*10^2}{50}=1000\:\rm N"